CF2Cl2 + 2Na2C2O4 = 2NaF + 2NaCl + C + 4CO2

What mass of NaCl in grams will be form if 2.35 mol of CF2Cl2 are heated and allowed to react completely with Na2C2O4?
1
Expert's answer
2020-11-20T08:07:34-0500

n(NaCl) = 2n(CF2Cl2)

n(NaCl) "= 2 \times 2.35 = 4.7 \;mol"

M(NaCl) = 58.44 g/mol

"m = n \times M"

m(NaCl) "= 4.7 \times 58.44 = 272.67 \;g"